# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ 4.2 Intrinsic Evaluation
4.3 Commonsense Reasoning
Table 6: 0-shot results (ACC) on commonsense reasoning: XCOPA, XStoryCloze, and XWinograd. EMMA-500 Llama 2 7B has better average performance than Llama 2 models and multilingual LLMs on XCOPA, XStoryCloze, and comparable performance on XWinograd.
Model XCOPA XStoryCloze XWinograd
Avg High Medium Low Avg High Medium Low Avg
Llama 2 7B 0.5667 0.6210 0.5390 0.5160 0.5755 0.6338 0.5445 0.4921 0.7247
Llama 2 7B Chat 0.5585 0.6125 0.5313 0.5060 0.5841 0.6480 0.5477 0.4974 0.6945
CodeLlama 2 7B 0.5469 0.5870 0.5253 0.5160 0.5568 0.6068 0.5233 0.4990 0.7092
LLaMAX Llama 2 7B 0.5438 0.5550 0.5413 0.5140 0.6036 0.6434 0.5882 0.5347 0.6749
LLaMAX Llama 2 7B Alpaca 0.5660 0.5980 0.5517 0.5240 0.6383 0.6908 0.6201 0.5433 0.6986
MaLA-500 Llama 2 10B v1 0.5309 0.5355 0.5327 0.5020 0.5307 0.5815 0.4922 0.4808 0.6589
MaLA-500 Llama 2 10B v2 0.5309 0.5355 0.5327 0.5020 0.5307 0.5815 0.4922 0.4808 0.6589
YaYi Llama 2 7B 0.5671 0.6210 0.5413 0.5060 0.5842 0.6498 0.5491 0.4904 0.7450
TowerBase Llama 2 7B 0.5633 0.6250 0.5290 0.5220 0.5778 0.6435 0.5367 0.4957 0.7429
TowerInstruct Llama 2 7B 0.5705 0.6290 0.5400 0.5200 0.5924 0.6683 0.5453 0.4970 0.7400
Occiglot Mistral 7B v0.1 0.5667 0.6280 0.5337 0.5200 0.5810 0.6518 0.5328 0.5003 0.7461
Occiglot Mistral 7B v0.1 Instruct 0.5655 0.6285 0.5297 0.5280 0.5939 0.6694 0.5433 0.5063 0.7293
BLOOM 7B 0.5689 0.5995 0.5587 0.5080 0.5930 0.6199 0.5905 0.5308 0.7013
BLOOMZ 7B 0.5487 0.5635 0.5460 0.5060 0.5712 0.6114 0.5582 0.4967 0.6795
mGPT 0.5504 0.5710 0.5440 0.5060 0.5443 0.5496 0.5453 0.5291 0.5969
mGPT 13B 0.5618 0.5975 0.5513 0.4820 0.5644 0.5776 0.5635 0.5331 0.6359
XGLM 7.5B 0.6064 0.6400 0.6037 0.4880 0.6075 0.6242 0.6036 0.5738 0.6884
YaYi 7B 0.5664 0.5955 0.5550 0.5180 0.6067 0.6490 0.5940 0.5265 0.6979
Llama 3 8B 0.6171 0.6835 0.5903 0.5120 0.6341 0.6850 0.6203 0.5344 0.7684
Llama 3.1 8B 0.6171 0.6930 0.5880 0.4880 0.6358 0.6866 0.6209 0.5387 0.7552
Gemma 7B 0.6364 0.7035 0.6143 0.5000 0.6501 0.6946 0.6449 0.5493 0.7741
Gemma 2 9B 0.6633 0.7340 0.6427 0.5040 0.6767 0.7247 0.6669 0.5764 0.8007
Qwen 2 7B 0.6031 0.6865 0.5640 0.5040 0.6146 0.6945 0.5697 0.5050 0.7644
Qwen 1.5 7B 0.5944 0.6685 0.5590 0.5100 0.5985 0.6662 0.5604 0.5056 0.7259
EMMA-500 Llama 2 7B 0.6311 0.6660 0.6257 0.5240 0.6638 0.6892 0.6573 0.6132 0.7280
@@ 4.4 Machine Translation
FLORES-200 is an evaluation benchmark for translation tasks with 204 language pairs involving English and thus 408 translation directions, with a particular focus on low-resource languages. We assess all language models by adopting a 3-shot evaluation approach with the prompt:
Translate the following sentence from {src_lang} to {tgt_lang}
[{src_lang}]: {src_sent}
[{tgt_lang}]:
The performance is measured by BLEU (98) and chrF++ (101) implemented in sacrebleu (102). The BLEU score is calculated with the flores200 tokenizer applied to the texts and chrF++ uses word order 2. The choice of flores200 tokenization ensures that languages that do not have a whitespace delimiter can be evaluated at the (sub-)word level. For reproducibility, we attach the BLEU and chrF++ signatures. [^19] BLEU: nrefs:1—case:mixed—eff:no—tok:flores200—smooth:exp—version:2.4.2,202020chrF++: nrefs:1—case:mixed—eff:yes—nc:6—nw:2—space:no—version:2.4.2
@@ 4.5 Text Classification
We use 3-shot prompting, drawing demonstrations from the development set and testing models on the test split. The prompt template for SIB-200 is as follows:
{examples}
The topic of the news "\${text}" is
For Taxi-1500, the prompt template is as follows:
{examples}
The topic of the verse "\${text}" is
@@
The outcomes are tabulated in Table 9 and Table 10 for SIB-200 and TAXI-1500 respectively. For SIB-200, our EMMA-500 model outperforms all Llama2-based models, with particularly notable gains in languages with medium or fewer resources—seeing an average improvement of 47.5%. Taxi-1500 could be a more challenging task since it is in the religious domain, but our model still surpasses all Llama2-based models except for MalA-500. However, despite these improvements in both classification tasks, our models lag behind the latest models such as Llama3 and 3.1, especially in high-resource languages.
@@
Table 9: 3-shot results on SIB-200 (ACC). EMMA-500 Llama 2 7B has better average performance than Llama 2 models and comparable performance with multilingual LLMs.
Model Avg High Medium-High Medium Medium-Low Low
Llama 2 7B 0.2241 0.2664 0.2469 0.2205 0.1968 0.1900
Llama 2 7B Chat 0.2558 0.2972 0.2811 0.2501 0.2303 0.2191
CodeLlama 2 7B 0.2335 0.2606 0.2542 0.2310 0.2142 0.2037
LLaMAX Llama 2 7B 0.1061 0.1242 0.1160 0.1001 0.0945 0.0954
LLaMAX Llama 2 7B Alpaca 0.2789 0.3309 0.3212 0.2716 0.2338 0.2282
MaLA-500 Llama 2 10B v1 0.2325 0.2330 0.2364 0.2288 0.2276 0.2358
MaLA-500 Llama 2 10B v2 0.1930 0.1893 0.2105 0.1949 0.1755 0.1846
Yayi Llama 2 7B 0.2457 0.2904 0.2717 0.2442 0.2144 0.2069
TowerBase Llama 2 7B 0.1934 0.2200 0.2092 0.1874 0.1790 0.1693
TowerInstruct Llama 2 7B 0.2053 0.2321 0.2196 0.2026 0.1915 0.1804
Occiglot Mistral 7B v0.1 0.3269 0.3880 0.3582 0.3174 0.2892 0.2836
Occiglot Mistral 7B v0.1 Instruct 0.3431 0.3948 0.3716 0.3336 0.3147 0.3008
BLOOM 7B 0.1781 0.2313 0.1805 0.1717 0.1576 0.1702
BLOOMZ 7B 0.2973 0.3039 0.2963 0.2980 0.2953 0.2970
mGPT 0.2711 0.2858 0.2799 0.2673 0.2589 0.2648
mGPT-13B 0.3320 0.3669 0.3427 0.3448 0.2939 0.3226
XGLM 7.5B 0.3181 0.3528 0.3512 0.3169 0.2696 0.2996
Yayi 7B 0.3576 0.4057 0.3620 0.3563 0.3472 0.3318
Llama 3 8B 0.6369 0.7345 0.7025 0.6462 0.5316 0.5696
Llama 3.1 8B 0.6142 0.7070 0.6790 0.6199 0.5146 0.5475
Gemma 7B 0.5821 0.6806 0.6455 0.5816 0.4886 0.5112
Gemma 2 9B 0.4625 0.5177 0.4900 0.4692 0.4304 0.4045
Qwen 1.5 7B 0.4795 0.5600 0.5181 0.4825 0.4156 0.4286
Qwen 2 7B 0.5495 0.6637 0.6014 0.5517 0.4519 0.4925
EMMA-500 Llama 2 7B 0.3127 0.3275 0.3328 0.3099 0.3083 0.2760
Table 10: 3-shot results on Taxi-1500 (ACC). EMMA-500 Llama 2 7B has comparable performance with the compared LLMs.
Model Avg High Medium-High Medium Medium-Low Low
Llama 2 7B 0.1754 0.1950 0.1949 0.1847 0.1746 0.1737
Llama 2 7B Chat 0.1544 0.1873 0.1766 0.1661 0.1559 0.1522
CodeLlama 2 7B 0.1703 0.1745 0.1741 0.1720 0.1705 0.1700
LLaMAX Llama 2 7B 0.2352 0.2320 0.2340 0.2376 0.2356 0.2352
LLaMAX Llama 2 7B Alpaca 0.1509 0.1870 0.1688 0.1599 0.1500 0.1491
MaLA-500 Llama 2 10B v1 0.2527 0.2390 0.2402 0.2476 0.2457 0.2543
MaLA-500 Llama 2 10B v2 0.2339 0.2136 0.2230 0.2132 0.2172 0.2367
Yayi Llama 2 7B 0.1773 0.1874 0.1846 0.1819 0.1789 0.1765
TowerBase Llama 2 7B 0.1773 0.1849 0.1881 0.1867 0.1810 0.1761
TowerInstruct Llama 2 7B 0.1729 0.2017 0.1960 0.1808 0.1740 0.1709
Occiglot Mistral 7B v0.1 0.2226 0.2464 0.2291 0.2299 0.2233 0.2215
Occiglot Mistral 7B v0.1 Instruct 0.1876 0.2430 0.2090 0.1941 0.1918 0.1848
BLOOM 7B 0.1476 0.1558 0.1489 0.1456 0.1511 0.1473
BLOOMZ 7B 0.1696 0.1693 0.1698 0.1696 0.1699 0.1695
mGPT 0.1072 0.0867 0.0844 0.0992 0.1029 0.1093
mGPT-13B 0.1723 0.1798 0.1644 0.1588 0.1610 0.1738
XGLM 7.5B 0.2041 0.2421 0.2369 0.2125 0.2105 0.2010
Yayi 7B 0.1612 0.1665 0.1638 0.1645 0.1583 0.1611
Llama 3 8B 0.2173 0.3184 0.2708 0.2560 0.2261 0.2105
Llama 3.1 8B 0.2020 0.2751 0.2521 0.2443 0.2097 0.1959
Gemma 2 9B 0.1805 0.2868 0.2855 0.2499 0.2028 0.1689
Gemma 7B 0.1383 0.2429 0.2413 0.1874 0.1497 0.1283
Qwen 1.5 7B 0.0729 0.1265 0.1145 0.0878 0.0730 0.0692
Qwen 2 7B 0.2187 0.2737 0.2557 0.2401 0.2233 0.2147
EMMA-500 Llama 2 7B 0.1982 0.2366 0.2333 0.2277 0.2200 0.1930
4.6 Open-Ended Generation
Aya Evaluation
Table 11: Results on Aya (BLEU/Self-BLEU). EMMA-500 Llama 2 7B has higher average BLEU scores than all baselines.
Model Avg High Medium-High Medium Medium-Low Low
Llama 2 7B 1.24/0.74 1.27/0.57 1.47/0.60 0.86/0.37 1.17/0.45 0.77/1.87
Llama 2 7B Chat 1.17/1.29 1.46/1.15 1.36/1.15 0.69/1.03 1.14/1.14 0.54/2.23
CodeLlama 2 7B 1.22/1.21 1.31/1.19 1.40/1.00 0.85/0.84 1.23/0.56 0.78/2.57
LLaMAX Llama 2 7B 1.72/1.70 1.80/1.37 2.03/1.48 1.30/1.24 1.65/0.89 0.98/3.74
LLaMAX Llama 2 7B Alpaca 1.68/1.67 1.82/1.22 1.96/1.42 1.28/1.14 1.55/1.04 1.00/3.94
MaLA-500 Llama 2 10B v1 0.40/2.29 0.42/2.53 0.49/2.79 0.32/1.22 0.31/2.42 0.18/1.16
MaLA-500 Llama 2 10B v2 0.41/2.31 0.42/2.01 0.50/2.65 0.32/1.42 0.33/1.02 0.18/3.09
Yayi Llama 2 7B 1.65/0.61 1.84/0.82 1.88/0.62 1.26/0.62 1.53/0.45 1.02/0.41
TowerBase Llama 2 7B 1.44/0.83 1.45/0.64 1.67/0.56 1.19/0.49 1.46/0.38 0.88/2.54
TowerInstruct Llama 2 7B 1.55/0.93 1.80/0.85 1.82/0.78 1.15/0.65 1.21/0.54 0.85/1.97
Occiglot Mistral 7B v0.1 1.53/2.43 1.57/0.91 1.78/2.89 1.17/1.73 1.61/1.31 0.95/4.27
Occiglot Mistral 7B v0.1 Instruct 0.75/2.81 0.82/2.38 0.89/3.10 0.43/2.72 0.79/1.17 0.45/3.50
BLOOM 7B 0.85/1.17 0.92/1.20 0.96/1.32 0.73/1.05 0.78/1.27 0.55/0.72
BLOOMZ 7B 0.12/0.61 0.07/0.33 0.17/0.62 0.08/1.00 0.06/0.89 0.08/0.51
mGPT 1.24/0.55 1.22/0.64 1.47/0.59 0.91/0.48 1.21/0.60 0.84/0.29
mGPT-13B 1.42/0.57 1.42/0.80 1.63/0.58 1.00/0.48 1.53/0.44 1.03/0.36
Yayi 7B 1.05/0.39 1.22/0.38 1.18/0.41 0.76/0.42 1.01/0.54 0.67/0.22
Llama 3 8B 1.59/0.94 1.03/0.60 1.31/0.53 2.96/0.54 1.11/0.28 2.43/3.47
Llama 3.1 8B 1.85/1.08 1.41/0.95 1.60/0.64 3.11/0.52 1.33/0.39 2.52/3.52
Gemma 2 9B 1.55/0.82 1.59/0.94 1.73/0.93 1.38/0.70 1.33/0.47 1.21/0.65
Gemma 7B 1.29/0.57 1.41/0.62 1.39/0.66 1.16/0.40 1.26/0.50 0.93/0.40
Qwen 1.5 7B 1.93/1.13 1.66/0.85 1.64/0.65 3.14/0.79 1.51/0.58 2.45/3.69
Qwen 2 7B 1.99/1.13 1.84/0.94 1.69/0.70 3.29/0.72 1.36/0.44 2.47/3.53
EMMA-500 Llama 2 7B 2.93/1.54 2.90/1.29 2.75/0.83 3.79/0.82 2.86/0.95 2.87/5.09
PolyWrite
This is a novel multilingual benchmark composed in this work for evaluating open-ended generation in 240 languages. We use ChatGPT to generate different prompts in English and use Google Translate to translate them into different languages for models to generate creative content. This benchmark consists of 31 writing tasks, such as storytelling and email writing, and 155 prompts in total. We back-translate the multilingual prompts to English, calculate the BLEU scores between original English prompts and back-translation, and filter out translated prompts with BLEU scores below 20, and the entire dataset contains a total of 35,751 prompts. We release the PolyWrite dataset on Huggingface [^22] https://huggingface.co/datasets/MaLA-LM/PolyWrite. The details of PolyWrite are described in Section C.1.
Table 12: Results on PolyWrite (Self-BLEU).
Model Avg High Medium-High Medium Medium-Low Low
Llama 2 7B 0.5358 0.4282 0.6545 0.3769 0.4766 0.6587
Llama 2 7B Chat 1.1550 0.8640 0.8902 1.1877 1.4435 1.4167
CodeLlama 2 7B 1.0313 1.2052 1.2883 0.9191 0.9092 0.6798
LLaMAX Llama 2 7B 0.9564 1.0066 1.1655 0.8786 0.8363 0.7709
LLaMAX Llama 2 7B Alpaca 0.8086 0.7321 0.9517 0.8322 0.8351 0.4981
MaLA-500 Llama 2 10B v1 3.7079 3.7902 3.5066 2.4541 4.8052 3.5163
MaLA-500 Llama 2 10B v2 4.0059 3.1737 4.0148 3.0476 5.0652 3.8916
Yayi Llama 2 7B 0.6274 0.6207 0.6921 0.4169 0.6813 0.6352
TowerBase Llama 2 7B 0.4938 0.7268 0.4736 0.3945 0.4417 0.5396
TowerInstruct Llama 2 7B 0.7124 0.9565 0.7651 0.6492 0.5998 0.6615
Occiglot Mistral 7B v0.1 0.9647 0.8818 0.7975 0.9543 0.9563 1.4157
Occiglot Mistral 7B v0.1 Instruct 3.9033 5.3884 4.7140 3.7555 2.8444 2.9568
BLOOM 7B 1.3892 1.2845 1.6705 1.9685 1.1513 0.6600
BLOOMZ 7B 0.0024 0.0000 0.0005 0.0093 0.0000 0.0049
mGPT 0.7560 0.9222 0.6291 0.4534 0.8156 1.1134
mGPT-13B 0.7479 0.8483 0.7091 0.4230 0.7962 1.0201
Yayi 7B 0.5151 0.4266 0.3574 0.3283 0.6706 0.8574
Llama 3 8B 0.5796 0.5753 0.5921 0.7141 0.5586 0.4440
Llama 3.1 8B 0.7995 0.6805 0.8253 0.5044 0.6965 1.3585
Gemma 2 9B 1.1736 1.2347 1.2913 1.1616 0.9261 1.3240
Gemma 7B 1.0541 0.9629 1.1222 0.9275 1.1112 1.0284
Qwen 1.5 7B 0.6441 0.9398 0.7457 0.3797 0.4164 0.8738
Qwen 2 7B 0.5709 0.7763 0.6135 0.3695 0.4186 0.7989
EMMA-500 Llama 2 7B 0.9879 0.9833 0.7058 0.8465 1.3130 1.1702

### v2

@@ 4.3 Machine Translation
FLORES-200 is an evaluation benchmark for translation tasks with 204 language pairs involving English and thus 408 translation directions, with a particular focus on low-resource languages. We assess all language models by adopting a 3-shot evaluation approach with the prompt in Section D.1.1. The performance is measured by BLEU (99) and chrF++ (102) implemented in sacrebleu (103). The BLEU score is calculated with the flores200 tokenizer applied to the texts and chrF++ uses word order 2. The choice of flores200 tokenization ensures that languages that do not have a whitespace delimiter can be evaluated at the (sub-)word level. For reproducibility, we attach the BLEU and chrF++ signatures. [^16] The BLEU signature is nrefs:1—case:mixed—eff:no—tok:flores200—smooth:exp—version:2.4.2; the chrF++ signature is nrefs:1—case:mixed—eff:yes—nc:6—nw:2—space:no—version:2.4.2
@@ 4.4 Text Classification
We use 3-shot prompting with prompts in Section D.1.2, drawing demonstrations from the development set and testing models on the test split. The outcomes are tabulated in Table 8 for SIB-200 and Taxi1500. For SIB-200, our EMMA-500 model outperforms all Llama2-based models, with particularly notable gains in languages with medium or fewer resources—seeing an average improvement of 47.5%. Taxi-1500 could be a more challenging task since it is in the religious domain, but our model still surpasses all Llama2-based models except for MalA-500. However, despite these improvements in both classification tasks, our models lag behind the latest models such as Llama3 and 3.1, especially in high-resource languages.
@@
Table 8: 3-shot results on SIB-200 and Taxi-1500 (ACC %). EMMA-500 Llama 2 7B has better average performance than Llama 2 models and comparable performance with multilingual LLMs, and has comparable performance with the compared LLMs.
Model SIB-200 Taxi-1500
Avg High Med-High Medium Med-Low Low Avg High Med-High Medium Med-Low Low
Llama 2 7B 22.41 26.64 24.69 22.05 19.68 19.00 17.54 19.50 19.49 18.47 17.46 17.37
Llama 2 7B Chat 25.58 29.72 28.11 25.01 23.03 21.91 15.44 18.73 17.66 16.61 15.59 15.22
CodeLlama 2 7B 23.35 26.06 25.42 23.10 21.42 20.37 17.03 17.45 17.41 17.20 17.05 17.00
LLaMAX Llama 2 7B 10.61 12.42 11.60 10.01 9.45 9.54 23.52 23.20 23.40 23.76 23.56 23.52
LLaMAX Llama 2 7B Alpaca 27.89 33.09 32.12 27.16 23.38 22.82 15.09 18.70 16.88 15.99 15.00 14.91
MaLA-500 Llama 2 10B v1 23.25 23.30 23.64 22.88 22.76 23.58 25.27 23.90 24.02 24.76 24.57 25.43
MaLA-500 Llama 2 10B v2 19.30 18.93 21.05 19.49 17.55 18.46 23.39 21.36 22.30 21.32 21.72 23.67
Yayi Llama 2 7B 24.57 29.04 27.17 24.42 21.44 20.69 17.73 18.74 18.46 18.19 17.89 17.65
TowerBase Llama 2 7B 19.34 22.00 20.92 18.74 17.90 16.93 17.73 18.49 18.81 18.67 18.10 17.61
TowerInstruct Llama 2 7B 20.53 23.21 21.96 20.26 19.15 18.04 17.29 20.17 19.60 18.08 17.40 17.09
Occiglot Mistral 7B v0.1 32.69 38.80 35.82 31.74 28.92 28.36 22.26 24.64 22.91 22.99 22.33 22.15
Occiglot Mistral 7B v0.1 Instruct 34.31 39.48 37.16 33.36 31.47 30.08 18.76 24.30 20.90 19.41 19.18 18.48
BLOOM 7B 17.81 23.13 18.05 17.17 15.76 17.02 14.76 15.58 14.89 14.56 15.11 14.73
BLOOMZ 7B 29.73 30.39 29.63 29.80 29.53 29.70 16.96 16.93 16.98 16.96 16.99 16.95
mGPT 27.11 28.58 27.99 26.73 25.89 26.48 10.72 8.67 8.44 9.92 10.29 10.93
mGPT-13B 33.20 36.69 34.27 34.48 29.39 32.26 17.23 17.98 16.44 15.88 16.10 17.38
Yayi 7B 35.76 40.57 36.20 35.63 34.72 33.18 16.12 16.65 16.38 16.45 15.83 16.11
Aya 23 8B 41.50 46.78 45.27 42.43 35.27 37.87 22.64 22.95 22.50 22.13 22.41 22.67
Aya expanse 8B 57.01 65.05 61.51 56.84 48.59 53.80 18.73 20.02 18.60 18.93 19.39 18.66
Llama 3 8B 63.69 73.45 70.25 64.62 53.16 56.96 21.73 31.84 27.08 25.60 22.61 21.05
Llama 3.1 8B 61.42 70.70 67.90 61.99 51.46 54.75 20.20 27.51 25.21 24.43 20.97 19.59
Gemma 7B 58.21 68.06 64.55 58.16 48.86 51.12 18.05 28.68 28.55 24.99 20.28 16.89
Gemma 2 9B 46.25 51.77 49.00 46.92 43.04 40.45 13.83 24.29 24.13 18.74 14.97 12.83
Qwen 1.5 7B 47.95 56.00 51.81 48.25 41.56 42.86 7.29 12.65 11.45 8.78 7.30 6.92
Qwen 2 7B 54.95 66.37 60.14 55.17 45.19 49.25 21.87 27.37 25.57 24.01 22.33 21.47
EMMA-500 Llama 2 7B 31.27 32.75 33.28 30.99 30.83 27.60 19.82 23.66 23.33 22.77 22.00 19.30
4.5 Commonsense Reasoning
Table 9: 0-shot results (ACC %) on commonsense reasoning: XCOPA, XStoryCloze, and XWinograd. EMMA-500 Llama 2 7B has better average performance than Llama 2 models and multilingual LLMs on XCOPA, XStoryCloze, and comparable performance on XWinograd.
Model XCOPA XStoryCloze XWinograd
Avg High Medium Low Avg High Medium Low Avg
Llama 2 7B 56.67 62.10 53.90 51.60 57.55 63.38 54.45 49.21 72.47
Llama 2 7B Chat 55.85 61.25 53.13 50.60 58.41 64.80 54.77 49.74 69.45
CodeLlama 2 7B 54.69 58.70 52.53 51.60 55.68 60.68 52.33 49.90 70.92
LLaMAX Llama 2 7B 54.38 55.50 54.13 51.40 60.36 64.34 58.82 53.47 67.49
LLaMAX Llama 2 7B Alpaca 56.60 59.80 55.17 52.40 63.83 69.08 62.01 54.33 69.86
MaLA-500 Llama 2 10B v1 53.09 53.55 53.27 50.20 53.07 58.15 49.22 48.08 65.89
MaLA-500 Llama 2 10B v2 53.09 53.55 53.27 50.20 53.07 58.15 49.22 48.08 65.89
YaYi Llama 2 7B 56.71 62.10 54.13 50.60 58.42 64.98 54.91 49.04 74.50
TowerBase Llama 2 7B 56.33 62.50 52.90 52.20 57.78 64.35 53.67 49.57 74.29
TowerInstruct Llama 2 7B 57.05 62.90 54.00 52.00 59.24 66.83 54.53 49.70 74.00
Occiglot Mistral 7B v0.1 56.67 62.80 53.37 52.00 58.10 65.18 53.28 50.03 74.61
Occiglot Mistral 7B v0.1 Instruct 56.55 62.85 52.97 52.80 59.39 66.94 54.33 50.63 72.93
BLOOM 7B 56.89 59.95 55.87 50.80 59.30 61.99 59.05 53.08 70.13
BLOOMZ 7B 54.87 56.35 54.60 50.60 57.12 61.14 55.82 49.67 67.95
mGPT 55.04 57.10 54.40 50.60 54.43 54.96 54.53 52.91 59.69
mGPT 13B 56.18 59.75 55.13 48.20 56.44 57.76 56.35 53.31 63.59
YaYi 7B 56.64 59.55 55.50 51.80 60.67 64.90 59.40 52.65 69.79
Aya 23 8B 55.13 59.05 53.30 50.40 60.93 67.27 58.82 49.34 65.84
Aya expanse 8B 56.38 61.50 53.77 51.60 64.80 73.10 61.93 49.80 71.94
Llama 3 8B 61.71 68.35 59.03 51.20 63.41 68.50 62.03 53.44 76.84
Llama 3.1 8B 61.71 69.30 58.80 48.80 63.58 68.66 62.09 53.87 75.52
Gemma 7B 63.64 70.35 61.43 50.00 65.01 69.46 64.49 54.93 77.41
Gemma 2 9B 66.33 73.40 64.27 50.40 67.67 72.47 66.69 57.64 80.07
Qwen 2 7B 60.31 68.65 56.40 50.40 61.46 69.45 56.97 50.50 76.44
Qwen 1.5 7B 59.44 66.85 55.90 51.00 59.85 66.62 56.04 50.56 72.59
EMMA-500 Llama 2 7B 63.11 66.60 62.57 52.40 66.38 68.92 65.73 61.32 72.80
4.6 Natural Language Inference
Table 10: 0-shot results on XNLI (ACC %).
Model Avg High Medium Low
Llama 2 7B 40.19 45.26 37.72 34.97
Llama 2 7B Chat 38.58 42.77 36.75 33.87
CodeLlama 2 7B 40.19 46.27 37.29 33.86
LLaMAX Llama 2 7B 44.27 46.53 42.64 43.03
LLaMAX Llama 2 7B Alpaca 45.09 48.47 42.80 42.89
MaLA-500 Llama 2 10B v1 38.11 42.10 35.85 34.65
MaLA-500 Llama 2 10B v2 38.11 42.10 35.85 34.65
YaYi Llama 2 7B 41.28 47.32 38.41 34.94
TowerBase Llama 2 7B 39.84 46.08 36.33 34.39
TowerInstruct Llama 2 7B 40.36 47.07 36.92 33.79
Occiglot Mistral 7B v0.1 42.35 49.90 38.39 35.19
Occiglot Mistral 7B v0.1 Instruct 40.81 47.58 37.18 34.52
BLOOM 7B 41.60 45.13 39.69 38.38
BLOOMZ 7B 37.13 40.02 35.56 34.51
mGPT 40.51 42.97 39.65 37.30
mGPT 13B 41.59 44.14 40.42 38.82
YaYi 7B 39.87 43.85 38.24 35.15
Aya 23 8B 43.12 48.51 41.95 34.67
Aya expanse 8B 45.56 50.48 44.24 38.38
Llama 3 8B 44.97 48.82 43.84 39.56
Llama 3.1 8B 45.62 49.61 44.04 40.83
Gemma 7B 42.58 46.44 41.00 38.01
Gemma 2 9B 46.74 48.50 45.11 46.49
Qwen 2 7B 42.77 47.31 41.35 36.53
Qwen 1.5 7B 39.47 40.95 38.80 37.83
EMMA-500 Llama 2 7B 45.14 46.09 44.40 44.71